# Resolving Task Objective Conflicts in Unified Multimodal Understanding and Generation via Task-Aware Mixture-of-Experts

### v1

@@ 4.1 Experimental Setups
| Qwen-VL-Chat Bai et al. (2023a) | 7B | - | 1487.5 | 60.6 | 58.2 | 57.5 | - | - | |
| Emu3-Chat Wang et al. (2024b) | 8B | 85.2 | 1244 | 58.5 | 68.2 | 60.3 | 31.6 | 37.2 | |
@@
| UTAMoE (Ours) | 3B | 87.2 | 1534.1 | 79.3 | 72.4 | 61.5 | 41.7 | 46.7 |
4.2 Quantitative Evaluation
Multimodal Understanding.
To assess multimodal understanding capabilities, we evaluate our model on widely recognized image-based vision-language benchmarks, which include POPE Li et al. (2023b), MME Fu et al. (2024), MMB Liu et al. (2024e), SEED Li et al. (2024a), GQA Hudson and Manning (2019), MMMU Yue et al. (2024) and MM-Vet Yu et al. (2023b).
Visual Generation.
For evaluating visual generation capabilities, we use GenEval Ghosh et al. (2023). GenEval is a challenging benchmark for text-to-image generation, designed to reflect the comprehensive generative abilities of visual generation models by offering a detailed instance-level analysis of their compositional capabilities.
Comparison Methods.
As shown in Table 2 and Table 3, we categorize the models used for comparison into three groups based on their capabilities. (i) Understanding Only: This group includes models designed solely for multimodal understanding, such as the LLaVA series (e.g., LLaVA Liu et al. (2023), LLaVA-v1.5 Liu et al. (2024c)), the MobileVLM series (e.g., MobileVLM Chu et al. (2023), MobileVLM-V2 Chu et al. (2024)), InstructBLIP Dai et al. (2023), Qwen-VL-Chat Bai et al. (2023a), and Emu3-Chat Wang et al. (2024b). (ii) Generation Only: This group contains models that focus exclusively on visual generation, including LlamaGen Sun et al. (2024), LDM Rombach et al. (2022), SDv1.5 Rombach et al. (2022), PixArt- Chen et al. (2023a), SDv2.1 Rombach et al. (2022), DALL·E 2 Ramesh et al. (2022), Emu3-Gen Wang et al. (2024b), SDXL Podell et al. (2023), DALL·E 3 Betker et al. (2023), and SD3-Medium Esser et al. (2024). (iii) Understanding and Generation: This group comprises models capable of both multimodal understanding and generation, such as DreamLLM Dong et al. (2023), LaViT Jin et al. (2023), MetaMorph Tong et al. (2024), NEXT-GPT Wu et al. (2024b), Show-o-256 Xie et al. (2024), SDHo-512 Xie et al. (2024), D-DiT Li et al. (2024c), Gemini-Nano-1 Team et al. (2023), ILLUME Wang et al. (2024a), TokenFlow-XL Qu et al. (2024), LWM Liu et al. (2024b), VILA-U Wu et al. (2024c), Chameleon Team (2024), and the Janus-series models Wu et al. (2024a); Chen et al. (2025); Ma et al. (2024).
@@ 4.2 Quantitative Evaluation
| UTAMoE (Ours) | 0.99 | 0.84 | 0.55 | 0.91 | 0.62 | 0.65 | 0.76 |
All in all, our UTAMoE, based on Chen et al. (2025), achieves state-of-the-art performance in multimodal understanding tasks, and delivers strong results in visual generation. This success is primarily attributed to our redesigned Mixture-of-Experts (MoE) mechanism, which effectively mitigates the task objective conflicts between multimodal understanding and generation, thereby significantly enhancing model’s overall capabilities.
4.3 Qualitative Evaluation
Figure 5 presents qualitative results for both multimodal understanding (left) and visual generation (right). UTAMoE demonstrates outstanding comprehension capabilities when processing inputs from diverse contexts, showcasing its strong capacity for multimodal semantic modeling, particularly in Basic Visual Description, Visual Question Answering (VQA), and Text Recognition tasks. The right side presents representative visual generation results across four categories: basic spatial relationships, natural landscapes, human characters, and imaginative scenes. Although the output resolution of UTAMoE is limited to 384 × 384, the model effectively captures the semantic intent of prompts, generating coherent, well-structured, and detail-rich images. These results highlight the superiority of our approach: UTAMoE not only models high-level semantic information in multimodal understanding tasks but also faithfully preserves fine-grained visual details in generation tasks, thereby mitigating inter-task objective conflicts.
4.4 Expert Load Analysis
Figure 4 illustrates expert utilization in understanding task, visualized by averaging values from randomly selected samples across multiple experimental runs. As shown in 4, Expert-0 and Expert-1 belong to the understanding group, with their combined activation rate exceeding 60% in most cases. Experts 2 and 3, belonging to the generation group, exhibit low activation rates across most layers, indicating that our task-aware router effectively routes tokens to the intended expert groups.Even when occasional routing errors increase generation group activations, the shared expert’s influence can partially compensate for meeting the task requirements. This further validates the effectiveness of our routing mechanism and shared expert design.
5 Ablation Study
In this section, we conduct a comprehensive ablation study to investigate the contributions of key components in our proposed framework. We focus on analyzing the effectiveness of Task-Aware MoE and the impact of our training strategy.
5.1 Effectiveness of Task-Aware MoE
@@
5.2 Effectiveness of Training Strategy
@@
6 Conclusion
In this paper, we propose a novel approach to the problem of Task Objective Conflicts in unified multimodal autoregressive models. The foundational idea of our approach is to decouple the internal components of autoregressive architecture. Specifically, we design UTAMoE, featuring a Task-Aware Mixture-of-Experts layer that creates task-specific optimization subpaths, effectively disentangling the conflicting objectives of understanding and generation tasks. Furthermore, we develop a Two-Stage Training Strategy that enhances task-specific differentiation while balancing objectives of distinct tasks. Extensive experimental results demonstrate that UTAMoE achieves state-of-the-art performance across various benchmarks.
7 Limitations
The diversity and inconsistency of current datasets, including variations in modality types, annotation quality, and task definitions, pose significant challenges to unified training and fair evaluation. Additionally, reliance on large pretrained backbone models results in substantial computational costs.
Appendix A Experimental Settings and Details
A.1 Details of Model Architecture
In our experiments, UTAMoE adopts a unified architecture for both multimodal understanding and generation tasks, inheriting the overall design from Chen et al. (2025). The architecture consists of decoupled vision encoder, aligners, text encoder, large language model (LLM), text decoder and image generation head. Except for LLM and decoupled vision encoder, all other components remain consistent across tasks.
Vision Encoder and Aligner.
For understanding tasks, we employ a vision encoder based on the CLIPVisionTower architecture with siglip-large-patch16-384 as the backbone, processing images at a resolution of 384×384. For generation tasks, we utilize a VQ-16 encoder with a downsampling factor of 16, producing 16,384 image tokens with a feature dimension of 8. Prior to inputting into the LLM, visual features from both encoders are projected into a unified embedding space via their respective alignment modules. Each alignment module is implemented as a two-layer MLP. the understanding aligner accepts inputs of dimension 1024 and outputs 2048-dimensional embeddings with GELU activation, while the generation aligner maps 8-dimensional inputs into the same 2048-dimensional space. The aligned features from both tasks are subsequently integrated into language model for subsequent multimodal fusion and processing.
Core Language Model.
Our language model is based on DeepSeek-LLM, supporting a maximum sequence length of 4,096 tokens. Building upon this foundation, we incorporate a Task-Aware Mixture-of-Experts (MoE) Layer to replace the original Transformer’s feed-forward network (FFN) modules, thereby resolving task objective conflict. Specifically, the model comprises 24 Transformer layers, each with the FFN replaced by our designed Task-Aware MoE layer containing two group-specific experts and one shared expert. The model has a hidden dimension of 2,048 and an intermediate dimension of 5,632, and is trained using bfloat16 precision. The vocabulary size is 102,400, and the maximum positional embedding supports sequence lengths up to 16,384 tokens. The multi-head attention mechanism is compatible with the LLaMA architecture.
Vision Head.
For the text-to-image generation task, we incorporate a vision head module to transform the language model outputs into discrete image tokens. This module is configured with an embedding dimension of 2048 and generates a total of 16,384 image tokens, enabling high-fidelity image reconstruction and visualization.
A.2 Training Dataset Details
We employ ShareGPT4V to train understanding tasks, while MidJourney is utilized for training generation tasks. Detailed descriptions of both datasets are provided below.
ShareGPT4V is a carefully curated collection of 102,000 high-quality image-text pairs, gathered from diverse sources including LAION, Conceptual Captions (CC), SBU Captions, SAM, COCO, web-landmark, web-celebrity, and Wikiart. Each pair was selected and refined to ensure accuracy, relevance, and quality. The GPT-4 Vision model was used during curation to analyze and improve captions, ensuring they align well with the images and are informative. This dataset is a valuable resource for researchers in computer vision and natural language processing, supporting tasks like image captioning, visual question answering, multimodal learning, content-based image retrieval, and art or landmark recognition. The ShareGPT4V dataset is stored in JSON format (see A.2), with each entry following a VQA format that includes image paths, questions, and answers—making it easy to load and process for model training.
MidJourney provides over 20 million prompt-image pairs encompassing diverse styles, compositions, and model versions, making it a valuable resource for training and research in text-to-image generation models. The dataset includes high-quality images paired with their corresponding prompts, as illustrated in A.2. In this work, we utilize a carefully selected subset of this dataset.
A.3 Evaluation Benchmarks
A.3.1 Multimodal Understanding
POPE: Evaluates object hallucination by converting hallucination detection into binary yes/no questions about object presence in MSCOCO images. It supports both annotated and raw images using automatic segmentation tools. Analysis shows 9% ambiguous and 5.5% incorrect annotations in the dataset.
MME: Covers 14 sub-tasks across perception, reasoning, and language understanding. It features diverse question types, instruction styles, and metrics, providing comprehensive diagnostic insights into MLLM performance.
MMBench: A multiple-choice evaluation benchmark in English and Chinese. It uses a Circular Evaluation strategy—shifting answer positions across multiple queries—to reduce bias and assess model consistency under varied choice orderings.
SEED-Bench: Tests interleaved vision-language comprehension and generation across three difficulty levels. Tasks include image-text understanding and multimodal instruction following, emphasizing compositional reasoning and complex visual input handling.
GQA: Focuses on compositional reasoning in visual question answering. Questions are generated from scene graphs to assess understanding of spatial relationships, attributes, and object roles in real-world images.
MMMU: Contains 11.5K multimodal questions drawn from textbooks and exams across six academic disciplines. Evaluates subject-specific reasoning with both visual and textual input, simulating high-level, expert-domain tasks.
MM-Vet: Assesses 16 combinations of six core skills: object recognition, commonsense knowledge, spatial understanding, text generation, OCR, and visual math. Uses LLM-based metrics for consistent scoring of open-ended answers.
A.3.2 Visual Generation
GenEval: Includes diverse challenges such as single-object rendering, object pairs, counting, spatial relations, and color accuracy. Evaluations are typically conducted using object detection models to assess generation fidelity and alignment with prompts.
DPG-Bench: Features prompts from multiple sources with an average length of 67 words, offering significant compositional and linguistic complexity. Generated images are evaluated via visual question answering models to test semantic alignment and content accuracy.
A.4 Image Processing: Any-resolution Strategy
Current unified models for multimodal understanding and generation typically employ fixed-resolution visual encoders. For instance, JanusPro utilizes the SigLIP Zhai et al. (2023) with a fixed 384384 input resolution for multimodal understanding tasks, which inevitably degrades performance on fine-grained tasks like text recognition due to loss of detailed information. Similarly, in visual generation, such resolution constraints lead to insufficient detail preservation in synthesized images.
Previous efforts Bai et al. (2023b); Li et al. (2023a) to enhance resolution often rely on positional embedding interpolation followed by fine-tuning Vision Transformer (ViT) backbone on large-scale image-text datasets. However, this approach not only demands substantial computational resources but also rigidly fixes the acceptable input resolution during inference. To overcome these challenges, we propose Any-Resolution strategy based Liu et al. (2024d) as illustrated in Figure 6.
Given an input image , we begin by dividing into local image patches.
Then, for each image patch , we randomly select one transformation from a set of predefined geometric transformations (e.g., rotation, shear, etc.) to apply to the patch.
Each transformed image is then encoded by visual encoder to produce visual tokens:
Simultaneously, the resized original image is encoded as:
All encoded features are flattened and concatenated to form a comprehensive representation:
Appendix B Supplementary Experiments
B.1 Validation of Task Objective Conflict
Comparative Analysis.
To investigate potential conflicts between task objectives, we conduct a comparative experiment between single-task and multi-task training. Specifically, we fine-tune two variants based on Chen et al. (2025) : (1) a single-task model, where each task is trained independently with its own objective function; and (2) a multi-task model, where all tasks are jointly optimized using a weighted combination of task-specific losses. We select POPE and Gen.Eval as representative benchmarks to cover both multimodal understanding and visual generation tasks. As shown in Table 6, if multi-task model shows noticeable performance drops on certain tasks and fails to achieve balanced performance across all tasks, this indicates the presence of inherent conflicts between task objectives.
| Task | POPE | GenEval | POPE | Gen.Eval | |
| (a) | Und. | 86.9 | – | – | – |
| (b) | Gen. | – | 0.74 | – | – |
| (c) | Both | 86.2 | 0.73 | -0.7 | -0.01 |
The results indicate the existence of task objective conflict. However, focusing solely on final performance metrics may overlook subtle interactions. To gain deeper insights, we further analyze the dynamics of the loss functions during training. Specifically, we examine the variations in the loss functions used for multimodal understanding and visual generation tasks. This detailed investigation helps confirm and understand the nature of these conflicts beyond just evaluation scores.
Loss Function Analysis.
To further verify the task objective conflict, we analyze the dynamics of the loss functions during multi-task training. Specifically, we track the loss values of multimodal understanding and visual generation tasks over training iterations to identify whether decreases in one task’s loss correspond to increases in another’s. Our understanding task uses cross-entropy loss, while the generation task employs mean squared error (MSE) loss. If significant mutual suppression in loss curves, it indicates the existence of optimization conflicts between tasks.
From Figure 7, we observe a certain degree of antagonism between the two loss functions, which further supports the existence of task conflicts. However, these experiments do not fundamentally verify the conflict, and addressing this at a deeper level remains a goal for our future work.
B.2 Generation Performance on DPG-Bench
DPG-Bench (Dense Prompt Graph Benchmark) comprises 1,065 complex and densely structured prompts, designed for evaluating visual generation capabilities, particularly in terms of fine-grained semantic alignment under linguistically and conceptually challenging scenarios. Results are summarized in Table 7.
| Method | Global | Entity | Attribute | Relation | Other | Overall |
| SDv1.5 Rombach et al. (2022) | 74.63 | 74.23 | 75.39 | 73.49 | 67.81 | 63.18 |
| PixArt- Chen et al. (2023a) | 74.97 | 79.32 | 78.60 | 82.57 | 76.96 | 71.11 |
| SDXL Podell et al. (2023) | 83.27 | 82.43 | 80.91 | 86.76 | 80.41 | 74.65 |
| Hunyuan-DiT Li et al. (2024b) | 84.59 | 80.59 | 88.01 | 74.36 | 86.41 | 78.87 |
| PixArt- Chen et al. (2024a) | 86.89 | 82.89 | 88.94 | 86.59 | 87.68 | 80.54 |
| Emu3-Gen Wang et al. (2024b) | 85.21 | 86.68 | 86.84 | 90.22 | 83.15 | 80.60 |
| DALL-E 3 Betker et al. (2023) | 90.97 | 89.61 | 88.39 | 90.58 | 89.83 | 83.50 |
| Janus Wu et al. (2024a) | 82.33 | 87.38 | 87.70 | 85.46 | 86.41 | 79.68 |
| JanusPro Chen et al. (2025) | 87.58 | 88.63 | 88.17 | 88.98 | 88.30 | 82.63 |
| UTAMoE (Ours) | 87.95 | 89.07 | 89.31 | 89.33 | 88.84 | 84.11 |
B.3 Comprehensive Analysis of Expert
The visualization analysis in subsection 4.4 illustrates the expert utilization patterns of our trained UTAMoE on the understanding task. However, several important questions remain to be explored. In the following subsection, we focus on addressing one key issue: What is the optimal ratio between group-specific experts and shared experts?
Optimal Ratio of Group-Specific and Shared Experts.
To explore the impact of the ratio between group-specific experts and shared experts on model performance, we conduct a series of experiments varying the G S ratio. Table 8 summarizes the results on POPE and GenEval benchmarks under different expert configurations. By analyzing these outcomes, we aim to identify the optimal balance that maximizes overall task performance and effectively leverages both specialized and shared knowledge within the model. Notably, we do not adopt the 3:1 ratio proposed by DeepSeek Dai et al. (2024), largely because we have redesigned the routing strategy, which alters the interaction between experts and affects the ideal ratio.
@@ B.3 Comprehensive Analysis of Expert
As presented in Table 8, our preliminary findings suggest that an optimal ratio of group-specific to shared experts is approximately 2. We emphasize that this conclusion is preliminary, and further extensive investigations will be conducted to validate and refine this observation in future work.
Appendix C Additional Visualizations
C.1 Multimodal Understanding.
In addition to the task types illustrated in Figure 5, several other representative tasks are widely studied in the field of multimodal understanding. Image-Text Matching focuses on determining whether a given image and a textual description semantically correspond to each other. Landmark Recognition focuses on detecting and classifying well-known landmarks or notable locations within an image. Visual Dialog requires engaging in a multi-turn conversation by jointly reasoning over both the visual content and the dialog history to generate coherent and contextually appropriate responses. To further highlight the robustness of our model’s multimodal understanding capabilities, we present additional qualitative examples in Figure 12.
C.2 Visual Generation.
Beyond scenarios presented in Figure 5, text-to-image generation also encompasses a wide range of additional contexts, including urban versus indoor, abstract versus concrete, and human-centered versus object-centered perspectives. The results are shown in Figure 13 below.

### v2

@@ 4.3. Quantitative Evaluation
| Qwen-VL-Chat Bai et al. (2023) | 7B | - | 1487.5 | 60.6 | 58.2 | 57.5 | - | - | |
| Emu3-Chat Wang et al. (2024c) | 8B | 85.2 | 1244 | 58.5 | 68.2 | 60.3 | 31.6 | 37.2 | |
@@
| UniDecouple (Ours) | 3B | 87.2 | 1534.1 | 79.3 | 72.4 | 61.5 | 41.7 | 46.7 |
| Method | Global | Entity | Attribute | Relation | Other | Overall |
| SDv1.5 Rombach et al. (2022) | 74.63 | 74.23 | 75.39 | 73.49 | 67.81 | 63.18 |
| PixArt- Chen et al. (2023b) | 74.97 | 79.32 | 78.60 | 82.57 | 76.96 | 71.11 |
| SDXL Podell et al. (2023) | 83.27 | 82.43 | 80.91 | 86.76 | 80.41 | 74.65 |
| Hunyuan-DiT Li et al. (2024c) | 84.59 | 80.59 | 88.01 | 74.36 | 86.41 | 78.87 |
| PixArt- Chen et al. (2024a) | 86.89 | 82.89 | 88.94 | 86.59 | 87.68 | 80.54 |
| Emu3-Gen Wang et al. (2024c) | 85.21 | 86.68 | 86.84 | 90.22 | 83.15 | 80.60 |
| DALL-E 3 Betker et al. (2023) | 90.97 | 89.61 | 88.39 | 90.58 | 89.83 | 83.50 |
| Janus Wu et al. (2024a) | 82.33 | 87.38 | 87.70 | 85.46 | 86.41 | 79.68 |
| JanusPro Chen et al. (2025) | 87.58 | 88.63 | 88.17 | 88.98 | 88.30 | 82.63 |
| UniDecouple (Ours) | 87.95 | 89.07 | 89.31 | 89.33 | 88.84 | 84.11 |
@@
| UniDecouple (Ours) | 0.99 | 0.84 | 0.55 | 0.91 | 0.62 | 0.65 | 0.76 |
All in all, our UniDecouple, based on Chen et al. (2025), preserves strong MMU ability while achieving T2I quality comparable to state-of-the-art methods. This success is primarily attributed to TA-MoE, which effectively mitigates TOC between MMU and T2I, thereby significantly enhancing model’s overall capabilities.
4.4. Qualitative Evaluation
Figure 6 presents qualitative results for both multimodal understanding (left) and text-to-image generation (right). UniDecouple demonstrates outstanding comprehension capabilities when processing inputs from diverse contexts, showcasing its strong capacity for multimodal semantic modeling, particularly in Basic Visual Description, Visual Question Answering (VQA), and Text Recognition tasks. The right side presents representative T2I results across four categories: basic spatial relationships, natural landscapes, human characters, and imaginative scenes. UniDecouple effectively captures the semantic intent of prompts, generating coherent, well-structured, and detail-rich images. These results highlight the superiority of our approach. UniDecouple not only models high-level semantic information in MMU tasks but also faithfully preserves fine-grained visual details in generation tasks, thereby mitigating inter-TOC.
expert-load
4.5. Expert Load Analysis
@@
Figure 5 illustrates expert utilization in understanding task, visualized by averaging values from randomly selected samples across multiple experimental runs. As shown in 5, Expert-0 and Expert-1 belong to the understanding group, with their combined activation rate exceeding 60% in most cases. Experts 2 and 3, belonging to the generation group, exhibit low activation rates across most layers, indicating that our task-aware router effectively routes tokens to the intended expert groups. Even when occasional routing errors increase generation group activations, the shared expert’s influence can partially compensate for meeting the task requirements. This further validates the effectiveness of our routing mechanism and shared expert design.
5. Ablation Study
In this section, we conduct a comprehensive ablation study to investigate the contributions of key components in our proposed framework. We focus on analyzing the effectiveness of Task-Aware MoE, the impact of training strategy and expert ratio.
5.1. Effectiveness of Task-Aware MoE
Vis
@@ 5.1. Effectiveness of Task-Aware MoE
5.2. Effectiveness of Training Strategy
@@
5.3. Impact of Expert Ratio
@@
The visualization in subsection 4.5 illustrates how UniDecouple utilizes experts on the understanding task, yet the optimal balance between group-specific and shared experts has not been fully established. To explore this, we conducted experiments on POPE and GenEval (Table 7), which suggest that a roughly 2:1 ratio yields the best results. Importantly, due to differences in routing strategy, we do not directly adopt the 3:1 ratio reported in DeepSeek Dai et al. (2024). This finding is preliminary and requires further investigation in future work.
6. Conclusion
In this paper, we address the challenge of Task Objective Conflicts in unified multimodal autoregressive models. We propose UniDecouple, a novel framework that decouples internal modules and constructs task-specific optimization subpaths, effectively mitigating negative transfer between multimodal understanding (MMU) and text-to-image generation (T2I) tasks. Central to our approach is the Task-Aware Mixture-of-Experts (TA-MoE), which combines Hierarchical Expert Routing and Hybrid Expert Collaboration, trained with a two-stage strategy to balance task specialization and overall coordination. Extensive experiments demonstrate that UniDecouple preserves strong understanding capabilities while achieving generation quality on par with state-of-the-art methods. Our results highlight the importance of disentangling task-specific pathways in unified models and provide a promising direction for future research on efficient, high-performing multimodal architectures.